OpenStudy (anonymous):

(1-cos)(1+cos)/cos^2 = tan^2 please help i would be so grateful

OpenStudy (anonymous):

(sin/1+sin)-(sin/1-sin)=-2tan^2

OpenStudy (phi):

is this the problem \[ \frac{\sin(x)}{1+\sin(x)} - \frac{\sin(x)}{1-\sin(x)}= -2 \tan^2(x)\]?

OpenStudy (phi):

I would multiply the two bottoms to find a common denominator.

OpenStudy (anonymous):

yes that is the problem

OpenStudy (anonymous):

f(x) = sin x/(1 + sin x) - sin x/(1 - sin x) = Common denominator: (1 - sin x)(1 + sin x) = (1 - sin^2 x) = cos ^2 x Numerator: sin x(1 - sin x) - sin x(1 + sin x) = = sin x - sin^2 x - sin x - sin^2 x = -2sin^2 x. f(x) = -2sin^2 x/(cos^2 x) = -2*tan^2 x The equation is proven.

OpenStudy (phi):

you should make it a new post... not edit an old one.

OpenStudy (anonymous):

i am new to this website

OpenStudy (phi):

ok, but please open a new post.

OpenStudy (anonymous):

can you help me

OpenStudy (phi):

Hit the "close button" on this post. and open a new one.